# A Comprehensive Survey on Uncertainty Quantification for Deep Learning

### v4

@@ 5.2.2. Deep Generative Model
VAE-based model: VAE model consists of two modules: an encoder and a decoder. The encoder network $q_{\phi}(\boldsymbol{z}|\boldsymbol{x})$ aims to embed the high dimensional structural output $\boldsymbol{x}$ into a low-dimensional code $\boldsymbol{z}$, that capture the inherent ambiguity or noise of the input data. The decoder $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$ aim to reconstruct the input feature. VAE model has been popular for structured output uncertainty, especially for the tasks on image data, because of their capability to model the global and local structure dependency in regular grid images. Specifically, two kinds of frameworks based on the VAE model have been proposed to account for the data uncertainty coming from input noise and target output noise: The first category aims to capture the noise that lies in the input samples. The basic idea is to embed each sample as a Gaussian distribution instead of a deterministic embedding in the low dimensional latent space, where the mean represents the feature embedding and variance represents the uncertainty of the embedding (Chang et al., 2020). The method takes into the different noise levels inherent in the dataset, which is ubiquitous in many kinds of real-world datasets, i.e., face image recognition (Chang et al., 2020), medical image reconstruction (Edupuganti et al., 2020). The probabilistic embedding framework takes advantage of the VAE model architecture to estimate the embedding and uncertainty simultaneously. The second category aims to capture the noise that lies in the target outputs, which also means the ground-truth is imperfect, ambiguous, or corrupted. This scenario is common in the medical domain (Lee et al., 2020), where the objects in the image are ambiguous and the experts may not reach a consensus on the class of the objects (large uncertainty). Thus for segmentation or classification tasks, the model should be aware of the prediction uncertainty. To capture the prediction uncertainty in the target outputs, the conditional VAE (cVAE) (Sohn et al., 2015)framework is adopted. Specifically, cVAE formulates the prediction distribution as an integration over the latent embedding $\boldsymbol{z}$,
@@
The cVAE model is trained by maximizing the evidence lower bound of the likelihood. Then during inference, multiple latent feature $\boldsymbol{z}_{j}$ can be drawn from the prior distribution and the integration over latent $\boldsymbol{z}$ can be approximated with the sampling distribution (Prokudin et al., 2018). Probabilistic U-Net model (Kohl et al., 2018) combines the architecture of cVAE and U-Net model by treating the U-Net model as the encoder to produce a probabilistic segmentation map. The U-Net model can capture multi-scale feature representations in the low-dimensional latent space to encode the potential variability in the segmentation map. These types of model are illustrated in Fig. 10 (a) and (b).
In summary, the VAE-based framework can take into consideration the data uncertainty coming from the input noise or the target output noise and can adapt the current state-of-the-art neural network architecture into its framework, making it more flexible for many kinds of applications. The key success lies in modeling the joint probability of all samples (pixels) in the image. The approach is suitable for structured uncertainty quantification (e.g., image grid structure, graph structure, et al.) through learning the implicit joint distribution on the structure.
@@
GAN-based generative model: GAN is a type of generative model trained with a two-player zero-game. It consists of a generator and a discriminator. In conditional GAN (cGAN), the generator takes the input $x$ and random noise $\boldsymbol{z}$ as input, and generates the target variables $y$: $\mathcal{G}:(\boldsymbol{x},\boldsymbol{z})\rightarrow y$. The discriminator is trained to distinguish the generated samples and ground-truth samples. The idea has been adopted in many domains. For example, in transportation domain, they adpot GAN-based approach for the prediction of traffic vollume (Mo and Fu, 2022). the flow model is integrated into GAN to enable likelihood estimation and better uncertainty quantification. Another approach (Oberdiek et al., 2022; Gao and Ng, 2022) further extends the method by utilizing the Wasserstein GAN (Arjovsky et al., 2017) based on gradient penalty to improve model convergence. The key advantage of deep generative modeling for uncertainty quantification is that they directly parameterize a deep neural network to represent the prediction distribution, without the need to have an explicit distribution format. Moreover, it can integrate a physics-informed neural network for better uncertainty estimation of physical science (Daw et al., 2021). However, the disadvantage is that deep generative models are much harder to train, especially for GAN-based models. The convergence of the model learning is not guaranteed.
@@ 5.3.1. Approaches combining data and model uncertainty
Combine BNN model with prediction distribution: The method aim to capture both data and model uncertainty in a single framework (Kendall and Gal, 2017) by combining the BNN and prediction distribution. The model uncertainty is captured with the BNN approximation approach. Specifically, MC drop-out is adopted due to its simplicity for implementation. For each dropout forward pass, one sample of the weight is drawn from the weight distribution approximation $\mathbf{W}_{t}\sim\text{Bernoulli}(p)$, where $p$ is the dropout rate, then one forward prediction can be made with the weight by $y_{t}=p(y|\boldsymbol{x},\mathbf{W}_{t})$. To obtain the data uncertainty, the output is formulated as a parameterized Gaussian distribution instead of point estimation $[y_{t},\sigma_{t}^{2}]=p(y|\boldsymbol{x};\mathbf{W}_{t})$, where $y_{t}$ is the target variable mean prediction and $\sigma_{t}^{2}$ is the prediction variance for a single forward prediction. With multiple dropout forward passes, we have a set of $T$ prediction samples $\{y_{t},\sigma^{2}_{t}\}_{t=1}^{T}$. The predictive uncertainty in the combined model can be approximated with the law of total variance expressed as $\text{Var}(y)$ in Eq. 20. The intuition of this equation is that the total uncertainty comes from two parts, the last $\frac{1}{T}\sum_{t=1}^{T}\sigma_{t}^{2}$ represents the average data uncertainty, and the first part $\frac{1}{T}\sum_{t=1}^{T}y_{t}^{2}-(\frac{1}{T}\sum_{t=1}^{T}y_{t})^{2}$ represent the disagreement of $T$ MC-dropout models, which capture the model uncertainty.
@@
Combine ensemble model with prediction distribution: This approach (Lakshminarayanan et al., 2017) combines the ensemble method with prediction distribution. The deep ensemble method constructs an ensemble of DNN models =$\mathcal{M}=\{\mathcal{M}_{i}\}_{i=1}^{K}$, where each model $\mathcal{M}_{i}$ can be set with different parameters, or architectures choices, et, al. The model uncertainty is expressed as the variance or ”disagreement” of the ensemble. In this way, the output of each model is modified as parameterized distribution to capture the data uncertainty. Similar to MC-dropout, we have an ensemble of prediction distribution $\{p(y|\boldsymbol{x},\mathcal{M}_{i})\}_{i=1}^{K}$. In this part, we take the classification problem as an example, where the prediction distribution is parameterized categorical distribution. The total uncertainty is captured with the entropy of average prediction distribution $\mathcal{H}(\mathbb{E}_{p(\mathcal{M}_{i})}p(y|\boldsymbol{x},\mathcal{M}_{i}))$, and the data uncertainty is the average entropy of each model, expressed as $\mathbb{E}_{p(\mathcal{M}_{i})}\mathcal{H}(p(y|\boldsymbol{x},\mathcal{M}_{i}))$. The model uncertainty can be expressed with the mutual information between the prediction and the ensemble model $y,\mathcal{M}$ as expressed in Eq. 21.
@@
In conclusion, to capture both the data and model uncertainty, existing literature can combine the methodologies in the two categories. There are several limitations to the combination approaches: first, the BNN or ensemble models require multiple forward passes for the prediction, which introduces computation overhead and extra storage. Efficiency is a concern. Second, the simple combination of data and model uncertainty lack a theoretical guarantee, which requires post-hoc calibration on the model.
@@ 5.3.2. Evidential deep learning
To overcome the computational challenge for the combination approaches, evidential deep learning was proposed to use one single deterministic model to capture both the data and model uncertainty without multiple forward passes of the neural network (Malinin and Gales, 2018; Charpentier et al., 2020; Sensoy et al., 2018; Bao et al., 2021). The intuition of evidential deep learning is to predict class-wise evidence instead of directly predicting class probabilities. In the following we review the methodologies, the advantages and disadvantages of those methods.
@@
Mathematically, evidential deep learning aims to learn the prior distribution of categorical distribution parameters, which is represented by the Dirichlet distribution. The Dirichlet distribution is parameterized by its concentration parameters $\boldsymbol{\alpha}$ (evidence) where $\alpha_{0}$ is the sum of all $\alpha_{i}$ and is referred to as precision of Dirichlet distribution. A higher value of $\alpha_{0}$ will lead to sharper distribution and lower model uncertainty. As shown in Eq.24, the $\text{Dir}(\boldsymbol{\mu}|\boldsymbol{\alpha})$ defines a probability density function over the k-dimensional random variable $\boldsymbol{\mu}=[{\mu_{1}},...,{\mu}_{k}]$, where $k$ is the number of classes, $\boldsymbol{\mu}$ belongs to the standard $k-1$ simplex (${\mu_{1}}+...+{\mu}_{k}=1$ and ${\mu}_{i}\in[0,1]$ for all $i\in{1,...,k}$), and can be regarded as the categorical distribution parameters. The relationship between Dirichlet distribution and uncertainty quantification can be illustrated using a 2-simplex. The random variable $\boldsymbol{\mu}=[{\mu}_{1},{\mu}_{2},{\mu}_{3}]$ is represented by its Barycentric coordinates in Fig 12 on the 2-simplex. The Barycentric coordinate is a coordinate system where point are located inside a simplex, and the value in each coordinate can be interpreted as the fraction of masses placed at the corresponding vertices of the simplex. Fig 12 (a) shows a scenario where the evidence parameters are equal, resulting in three classes being indistinguishable and implying high data uncertainty (high entropy for the sampled $\boldsymbol{\mu}$). As the total evidence $\sum_{i}\alpha_{i}$ increases, the density becomes more concentrated, which means the model uncertainty decreases, while the data uncertainty remains fixed. Fig 12 (b) shows a scenario with fixed model uncertainty, as the sum of evidence parameters remains constant. When the evidence becomes imbalanced, the density becomes more concentrated toward one class, thus decreasing the data uncertainty.
@@
(b) Keep the model uncertainty fixed, and the data uncertainty decrease from left to right (the entropy of the categoric distribution decrease).
@@
The advantage of evidential deep learning is the approach only requires a single forward pass during inference, and is much more computationally efficient. The approach also explicitly distinguishes data and model uncertainty in a principled way. The disadvantage is that the training stage is more complicated and is not guaranteed to obtain the same or similar prediction accuracy with vanilla network and cannot leverage existing progress in DNN. Furthermore, the training stage requires OOD samples to learn well representations and increase the amount of dataset.
@@
Computational efficient.
Performance degradation
Only capture OOD
uncertainty

### v5

@@ 5.2.2. Deep Generative Model
VAE-based model: VAE model consists of two modules: an encoder and a decoder. The encoder network $q_{\phi}(\boldsymbol{z}|\boldsymbol{x})$ aims to embed the high dimensional structural output $\boldsymbol{x}$ into a low-dimensional code $\boldsymbol{z}$, that capture the inherent ambiguity or noise of the input data. The decoder $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$ aim to reconstruct the input feature. VAE model has been popular for structured output uncertainty, especially for the tasks on image data, because of their capability to model the global and local structure dependency in regular grid images. Specifically, two kinds of frameworks based on the VAE model have been proposed to account for the data uncertainty coming from input noise and target output noise: The first category aims to capture the noise that lies in the input samples. The basic idea is to embed each sample as a Gaussian distribution instead of a deterministic embedding in the low dimensional latent space, where the mean represents the feature embedding and variance represents the uncertainty of the embedding (Chang et al., 2020). The method takes into the different noise levels inherent in the dataset, which is ubiquitous in many kinds of real-world datasets, i.e., face image recognition (Chang et al., 2020), medical image reconstruction (Edupuganti et al., 2020). The probabilistic embedding framework takes advantage of the VAE model architecture to estimate the embedding and uncertainty simultaneously. The second category aims to capture the noise that lies in the target outputs, which also means the ground truth is imperfect, ambiguous, or corrupted. This scenario is common in the medical domain (Lee et al., 2020), where the objects in the image are ambiguous and the experts may not reach a consensus on the class of the objects (large uncertainty). Thus for segmentation or classification tasks, the model should be aware of the prediction uncertainty. To capture the prediction uncertainty in the target outputs, the conditional VAE (cVAE) (Sohn et al., 2015)framework is adopted. Specifically, cVAE formulates the prediction distribution as an integration over the latent embedding $\boldsymbol{z}$,
@@
The cVAE model is trained by maximizing the evidence lower bound of the likelihood. Then during inference, multiple latent feature $\boldsymbol{z}_{j}$ can be drawn from the prior distribution, and the integration over latent $\boldsymbol{z}$ can be approximated with the sampling distribution (Prokudin et al., 2018). Probabilistic U-Net model (Kohl et al., 2018) combines the architecture of cVAE and U-Net model by treating the U-Net model as the encoder to produce a probabilistic segmentation map. The U-Net model can capture multi-scale feature representations in the low-dimensional latent space to encode the potential variability in the segmentation map. These types of models are illustrated in Fig. 11 (a) and (b).
In summary, the VAE-based framework can take into consideration the data uncertainty coming from the input noise or the target output noise and can adapt the current state-of-the-art neural network architecture into its framework, making it more flexible for many kinds of applications. The key success lies in modeling the joint probability of all samples (pixels) in the image. The approach is suitable for structured uncertainty quantification (e.g., image grid structure, graph structure, et al.) through learning the implicit joint distribution of the structure.
@@
GAN-based generative model: GAN is a type of generative model trained with a two-player zero-game. It consists of a generator and a discriminator. In conditional GAN (cGAN), the generator takes the input $x$ and random noise $\boldsymbol{z}$ as input and generates the target variables $y$: $\mathcal{G}:(\boldsymbol{x},\boldsymbol{z})\rightarrow y$. The discriminator is trained to distinguish the generated samples and ground-truth samples. The idea has been adopted in many domains. For example, in the transportation domain, they adpot a GAN-based approach for the prediction of traffic vollume (Mo and Fu, 2022). the flow model is integrated into GAN to enable likelihood estimation and better uncertainty quantification. Another approach (Oberdiek et al., 2022; Gao and Ng, 2022) further extends the method by utilizing the Wasserstein GAN (Arjovsky et al., 2017) based on gradient penalty to improve model convergence. The key advantage of deep generative modeling for uncertainty quantification is that they directly parameterize a deep neural network to represent the prediction distribution, without the need to have an explicit distribution format. Moreover, it can integrate a physics-informed neural network for better uncertainty estimation of physical science (Daw et al., 2021). However, the disadvantage is that deep generative models are much harder to train, especially for GAN-based models. The convergence of the model learning is not guaranteed.
@@ 5.3.1. Approaches combining data and model uncertainty
Combine BNN model with prediction distribution: The method aims to capture both data and model uncertainty in a single framework (Kendall and Gal, 2017) by combining the BNN and prediction distribution. The model uncertainty is captured with the BNN approximation approach. Specifically, MC drop-out is adopted due to its simplicity for implementation. For each dropout forward pass, one sample of the weight is drawn from the weight distribution approximation $\mathbf{W}_{t}\sim\text{Bernoulli}(p)$, where $p$ is the dropout rate, then one forward prediction can be made with the weight by $y_{t}=p(y|\boldsymbol{x},\mathbf{W}_{t})$. To obtain the data uncertainty, the output is formulated as a parameterized Gaussian distribution instead of point estimation $[y_{t},\sigma_{t}^{2}]=p(y|\boldsymbol{x};\mathbf{W}_{t})$, where $y_{t}$ is the target variable mean prediction and $\sigma_{t}^{2}$ is the prediction variance for a single forward prediction. With multiple dropout forward passes, we have a set of $T$ prediction samples $\{y_{t},\sigma^{2}_{t}\}_{t=1}^{T}$. The predictive uncertainty in the combined model can be approximated with the law of total variance expressed as $\text{Var}(y)$ in Eq. 19. The intuition of this equation is that the total uncertainty comes from two parts, the last $\frac{1}{T}\sum_{t=1}^{T}\sigma_{t}^{2}$ represents the average data uncertainty, and the first part $\frac{1}{T}\sum_{t=1}^{T}y_{t}^{2}-(\frac{1}{T}\sum_{t=1}^{T}y_{t})^{2}$ represent the disagreement of $T$ MC-dropout models, which capture the model uncertainty.
@@
Combine ensemble model with prediction distribution: This approach (Lakshminarayanan et al., 2017) combines the ensemble method with prediction distribution. The deep ensemble method constructs an ensemble of DNN models =$\mathcal{M}=\{\mathcal{M}_{i}\}_{i=1}^{K}$, where each model $\mathcal{M}_{i}$ can be set with different parameters, or architectures choices, et, al. The model uncertainty is expressed as the variance or "disagreement" of the ensemble. In this way, the output of each model is modified as a parameterized distribution to capture the data uncertainty. Similar to MC-dropout, we have an ensemble of prediction distribution $\{p(y|\boldsymbol{x},\mathcal{M}_{i})\}_{i=1}^{K}$. In this part, we take the classification problem as an example, where the prediction distribution is a parameterized categorical distribution. The total uncertainty is captured with the entropy of average prediction distribution $\mathcal{H}(\mathbb{E}_{p(\mathcal{M}_{i})}p(y|\boldsymbol{x},\mathcal{M}_{i}))$, and the data uncertainty is the average entropy of each model, expressed as $\mathbb{E}_{p(\mathcal{M}_{i})}\mathcal{H}(p(y|\boldsymbol{x},\mathcal{M}_{i}))$. The model uncertainty can be expressed with the mutual information between the prediction and the ensemble model $y,\mathcal{M}$ as expressed in Eq. 20.
@@
In conclusion, to capture both the data and model uncertainty, existing literature can combine the methodologies in the two categories. There are several limitations to the combination approaches: first, the BNN or ensemble models require multiple forward passes for the prediction, which introduces computation overhead and extra storage. Efficiency is a concern. Second, the simple combination of data and model uncertainty lacks a theoretical guarantee, which requires post hoc calibration on the model.
@@ 5.3.2. Evidential deep learning
To overcome the computational challenge for the combination approaches, evidential deep learning was proposed to use one single deterministic model to capture both the data and model uncertainty without multiple forward passes of the neural network (Malinin and Gales, 2018; Charpentier et al., 2020; Sensoy et al., 2018; Bao et al., 2021). The intuition of evidential deep learning is to predict class-wise evidence instead of directly predicting class probabilities. In the following section, we review the methodologies and their advantages and disadvantages.
@@
Mathematically, evidential deep learning aims to learn the prior distribution of categorical distribution parameters, which is represented by the Dirichlet distribution. The Dirichlet distribution is parameterized by its concentration parameters $\boldsymbol{\alpha}$ (evidence) where $\alpha_{0}$ is the sum of all $\alpha_{i}$ and is referred to as the precision of Dirichlet distribution. A higher value of $\alpha_{0}$ will lead to sharper distribution and lower model uncertainty. As shown in Eq.23, the $\text{Dir}(\boldsymbol{\mu}|\boldsymbol{\alpha})$ defines a probability density function over the k-dimensional random variable $\boldsymbol{\mu}=[{\mu_{1}},...,{\mu}_{k}]$, where $k$ is the number of classes, $\boldsymbol{\mu}$ belongs to the standard $k-1$ simplex (${\mu_{1}}+...+{\mu}_{k}=1$ and ${\mu}_{i}\in[0,1]$ for all $i\in{1,...,k}$), and can be regarded as the categorical distribution parameters. The relationship between Dirichlet distribution and uncertainty quantification can be illustrated using a 2-simplex. The random variable $\boldsymbol{\mu}=[{\mu}_{1},{\mu}_{2},{\mu}_{3}]$ is represented by its Barycentric coordinates in Fig 13 on the 2-simplex. The Barycentric coordinate is a coordinate system where points are located inside a simplex, and the value in each coordinate can be interpreted as the fraction of masses placed at the corresponding vertices of the simplex. Fig 13 (a) shows a scenario where the evidence parameters are equal, resulting in three classes being indistinguishable and implying high data uncertainty (high entropy for the sampled $\boldsymbol{\mu}$). As the total evidence $\sum_{i}\alpha_{i}$ increases, the density becomes more concentrated, which means the model uncertainty decreases, while the data uncertainty remains fixed. Fig 13 (b) shows a scenario with fixed model uncertainty, as the sum of evidence parameters remains constant. When the evidence becomes imbalanced, the density becomes more concentrated toward one class, thus decreasing the data uncertainty.
@@
(b) Keep the model uncertainty fixed, and the data uncertainty decreases from left to right (the entropy of the categoric distribution decreases).
@@
The advantage of evidential deep learning is the approach only requires a single forward pass during inference and is much more computationally efficient. The approach also explicitly distinguishes data and model uncertainty in a principled way. The disadvantage is that the training stage is more complicated and is not guaranteed to obtain the same or similar prediction accuracy with vanilla network and cannot leverage existing progress in DNN. Furthermore, the training stage requires OOD samples to learn well representations and increase the amount of dataset.
@@
Capture both data and
model uncertainty and
Computationally efficient
compared with BNN and
ensemble methods
Strong assumption on prior distribution. Difficult to train compared with BNN or ensemble.
Conformal
Prediction
Prediction set
or interval (Lu et al., 2022; Caprio et al., 2023; Hechtlinger et al., 2018; Karimi and Samavi, 2023)
Capture both data and
model uncertainty and
Distribution free
Strong assumption on sample distribution.
5.3.3. Conformal Prediction
Though existing UQ approaches can quantify total uncertainty from different sources, the lack of formal guarantees is a major limitation in their methods. To address this issue, conformal prediction (CP) is a post-process approach by constructing a finite prediction set with the statistical guarantee to cover the true label (Angelopoulos et al., 2023).
Conformal prediction is a method of distribution-free uncertainty quantification, which is model-agnostic and provides formal results for marginal coverage, defined as the the average probability that the true class will be contained in the prediction set (Angelopoulos and Bates, 2021). Formally, this coverage guarantee states:
(25) $\mathbb{P}(y\in\mathcal{C}(x))>=1-\alpha$
where $\boldsymbol{x}$ is a sample instance, and $y$ is the ground truth. Unlike traditional confidence intervals which typically assume a specific distribution of data or residuals, conformal prediction offers a non-parametric and distribution-free approach (Karimi and Samavi, 2023). This is advantageous in deep learning, where the assumptions required by traditional methods are often violated due to the complexity and non-linearity of the models.
There are several approaches to developing conformal predictions for deep learning models. The first kind of approach is to apply the conformal method to pre-trained deep learning models directly (Lu et al., 2022). Second, credal Bayesian deep learning trains an (uncountable) infinite ensemble of BNNs using only finitely many elements, and outputs a prediction set for total uncertainty estimation (Caprio et al., 2023). Third, density-based deep conform prediction considers the distance among samples to model the uncertainty coming from out-of-distribution aspects explicitly (Hechtlinger et al., 2018). However, the limitation of conformal prediction in deep learning is that it depends on the assumption that the data instances are independently and identically distributed (i.i.d.) and exchangeable. This assumption is violated in many structured data, including spatiotemporal data, and graph data, where there exists structural dependency among samples.
5.3.4. Evaluation Metrics
The previous subsections discuss various approaches for model and data uncertainty quantification. In this part, we discuss how to evaluate the performance of an uncertainty quantification model. There are several metrics to evaluate the performance of uncertainty quantification from different perspectives. First, one method is to evaluate the calibration between the uncertainty measure and the fraction of correct prediction, called expected calibration error (ECE (Naeini et al., 2015)). It first discretizes the uncertainty value into a fixed number of bins and assigns each predicted probability to the bin that encompasses it. The calibration error is the difference between the fraction of predictions in the bin that are correct (accuracy) and the mean of the probabilities in the bin (confidence). Given the groundtruth $\hat{y}$, the prediction from a model $Y$, and prediction uncertainty $\hat{p}$, the model is calibrated when the probability of being correct $\mathbb{P}(Y=\hat{y})$ is equal to the uncertainty measure $\hat{p}$:
(26) $\mathbb{P}(Y=\hat{y}|\hat{p})=\hat{p}\ \text{for any}\ \hat{p}$
Second, to address any distortion of results introduced by the binning procedure, the smooth ECE (smECE (Błasiok and Nakkiran, 2023)), which avoids the binning altogether by smoothing observations using an RBF kernel. Third, the area under the ROC (AUROC) (Ovadia et al., 2019) treats the problem as a binary misclassification by using uncertainty to predict the correctness of the prediction and aggregating the results over all possible decision thresholds. The intuition of this measure is that more uncertainty should indicate a lower probability of the correctness of the answers. Thus, AUROC measures the rank of the uncertainty score. Fourth, the Brier score is a strictly proper score function (Brier, 1950). It measures the mean squared difference between the predicted probability and the actual outcome. Last, Adaptive Calibration Error (ACE (Nixon et al., 2019)) computes the calibration across all predictions for multi-classes to overcome the limitation of ECE.